# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandit Environments

### v2

@@ 3.1 Synthetic Experiments
Figure 1: Dynamic regret vs. environment steps for synthetic experiments. First three columns correspond to the geometric change-points and the final one to the drifting case.
@@ 3.1.1 Experimental Parameters
In both NS-PB and NS-KB synthetic experiments, certain parameters are chosen to be the same. Specifically, in all synthetic experiments, the action space comprises $100$ unique actions with dimension $d=10$. These actions are sampled independently from $\mathcal{N}(0,I)$. The horizon is fixed to $T=50000$ and we average the results over $15$ independent runs.
@@
In Algorithm 1, allowing the change detector to use samples from the stationary bandit algorithm $\mathcal{B}$ when $|\mathcal{A}|$ is finite, but only allowing forced exploration over actions in $\mathcal{A}_{\mathrm{e}}$ improves empirical performance. This variation does not affect the theoretical properties of the algorithm, and we employ this variation for our experiments.
@@
The actions are scaled to lie within an $L$-ball and the underlying parameters $\theta_{t}$ belong to an $S$-ball. Specifically, for NS-LBs and NS-GLBs we have that $S=L=1$, while for NS-SCBs, we have that $L=1$ but $S=3$. Every time a $\theta_{t}$ is initialized or changed, its elements are chosen independently and uniformly in $[-1,1]$, and then are scaled to the $S$-ball. For both GLB and SCB, we select $\mu(x):=(1+e^{-x})^{-1}$. The additive noise is sampled according to $\mathcal{N}(0,0.01)$ at each time-step, while for the SCB setting, we sample the random reward according to $\mathrm{Bernoulli}(\mu(\langle\theta_{t},A_{t}\rangle))$ at time $t$. To select $\mathcal{A}_{\mathrm{e}}$ in the NS-PB setting we use Corollary 4.4.
@@
Actions are scaled in the $\sqrt{d}$-ball. Regarding the choice of kernel for NS-KBs, we employ the SE kernel with parameter $\ell=0.2$. The noise is sampled according to $\mathcal{N}(0,0.01)$ at each time-step. In the NS-KB setting, we follow a procedure similar to [14, 17]. Specifically, every time we initialize or change the reward function, $f_{t}$ is generated from the RKHS obtained by a discretization of $[-1,1]$ into 200 evenly spaced points. The reward function is generated as $f(\cdot)=\sum_{i=1}^{M}\alpha_{i}k(\cdot,x_{i})$ with $\alpha_{i}$ uniformly distributed in $[-1,1]$ and $M=200$. To identify $\mathcal{A}_{\mathrm{e}}$, we use Corollary 4.4.
@@ 3.1.2 Experimental Benchmarks
To compare the algorithms in the PS setting, we adopt the geometric change-point model proposed in [25], and independently sample the intervals between the change-points according to a geometric distribution with parameter $\rho=T^{-\xi}$, for $\xi\in\{0.4,0.6,0.8\}$. The expected number of change-points then becomes $\rho T=T^{1-\xi}$, which is sublinear in $T$ for chosen values of $\xi$. Our experiments do not have any restriction on the lengths of the intervals between change-points.
@@
Regarding comparisons in drifting non-stationarity, we adopt the following drift model: in each run, the reward structure changes linearly over $T$ rounds from an initial value to a final value, where the end-points are chosen the same way described in the beginning of the section. Specifically, we have that
[…]
@@
Per the results in Figure 1, DAL outperforms the current state-of-art methods in every synthetic experiment. The only competitive method at certain scenarios is WeightUCB, which, unlike DAL, is prior-based. Notably, DAL’s efficacy against the PS benchmarks significantly increases with the decrease in the number of change-points. DAL only abandons the action choices of the stationary bandit algorithm and restarts learning when an efficient change detector flags a mean-shift in rewards; hence, it avoids unnecessary restarts, especially when the intervals between the change-points are long enough for such detectors to correctly flag said changes without false alarms. We suspect that the judicious nature of the restarts is behind its observed superior performance, when changes are further apart. Regarding drifting non-stationarity, DAL significantly outperforms all other methods. In fact, it fares better than both WeightUCB and ADA-OPKB, which not only are known to attain the optimal regret bound in the drift setup, but have also been shown to perform well in practice.
@@ 3.2 Real-World Experiments
Figure 2: Results for real-world experiments of Section 3.2 averaged over 15 independent runs.
Click Log Benchmark.
We construct an NS bandit benchmark from the Yahoo! R6A click log dataset. [^1] Yahoo! Front Page Today Module User Click Log Dataset: https://webscope.sandbox.yahoo.com. Following prior works [9, 50], we compute smoothed click-through rates (CTRs) via centered rolling averages over 2000 rounds, average CTRs within each subperiod, and suppress fluctuations below $0.005$. To increase difficulty, we combine actions across 5 days, leading to 64 actions, compress the horizon to $50000$, and multiply final CTRs by 10. We model the resulting environment as an NS-SCB problem, reflecting the logistic reward structure typical in such settings [44].
@@
NS-KBs have been applied to stock market prediction, and we follow the procedure of [17] to simulate two environments: one using their original data (Quandl stocks) and one constructed from NASDAQ-100 stocks retrieved via the yfinance Python package. [^2] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes. In the Yahoo-based dataset, we retain stocks with sufficient history ($T{=}2000$, approx. 5.5 years) and select the $K=50$ most volatile as actions. Daily closing prices define the reward function, and the empirical price covariance matrix is used as the kernel. To increase difficulty, we add Gaussian noise $\mathcal{N}(0,0.01)$ to the reward at each time-step.
@@
We construct a NS bandit environment based on the Criteo live traffic dataset [18], following the preprocessing approach of [45] but modeling the problem as an NS-GLB rather than an NS-LB. We estimate the underlying parameter $\theta^{*}$ using logistic regression. Unlike [45], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$ and extend the horizon to $T=50000$.
@@
We use the Bioliq dataset provided by [31], which contains a week of measurements from 20 sensors in a power plant. We process the reward as [31] and construct an NS-SCB environment with $190$ actions. At each time-step, the reward is $1$ if the last 1000 measurements exceed a threshold of $2.04$, and 00 otherwise. Evaluation is based on cumulative reward.
@@ 4.1 On Effective Detection
Effective non-stationarity detection is crucial for strong practical performance. When selecting a non-stationarity detector, it is essential to ensure accuracy, efficiency, and optimal regret growth properties. Any detector aiming to identify distribution shifts inherently requires a certain number of samples, both before and after the change occurs. Ideally, this sample complexity should scale appropriately to avoid negatively impacting the final regret. To this end, it has been demonstrated that the GLR test achieves a pre- and post-change sample complexity that scales as $\mathcal{O}(\log T)$ [28]. Because logarithmic terms are disregarded in dynamic regret analyses, integrating this detection mechanism properly suggests that optimal regret growth can be achieved.
However, in order to select which samples should be fed into the change detector, one needs to properly design the covering set $\mathcal{A}_{\mathrm{e}}$. When designing $\mathcal{A}_{\mathrm{e}}$, one needs to ensure that it contains actions that can capture the changes in the reward function. However, such changes cannot be allowed to be made arbitrarily small, as no change detector may be able to identify them. $\mathcal{A}_{\mathrm{e}}$ should be designed such that whenever an arbitrary change occurs, reward sequences associated with at least one action in $\mathcal{A}_{\mathrm{e}}$ exhibit an appreciable mean-shift. Define
$\Delta_{\mathrm{c}}\coloneqq\inf_{f\neq f^{\prime}}\max_{a\in\mathcal{A}_{\mathrm{e}}}\ \lvert f(a)-f^{\prime}(a)\rvert.$
Then, $\Delta_{\mathrm{c}}$ captures how well the actions in $\mathcal{A}_{\mathrm{e}}$ can discern between candidate reward functions. According to [27], $\Delta_{\mathrm{c}}$ crucially affects the performance of the GLR test, as its pre- and post- change sample complexity grows with $1/\Delta_{\mathrm{c}}^{2}$. The more discernible the changes are, the easier the detection becomes. Since forced exploration incurs regret, $\mathcal{A}_{\mathrm{e}}$ should ideally be chosen to minimize $N_{\mathrm{e}}$ while maximizing $\Delta_{\mathrm{c}}$. However, this cannot be done since the function $f_{t}$ is unknown. Hence, we provide the ways with which one can ensure that $\mathcal{A}_{\mathrm{e}}$ can lead to detecting the changes in $f_{t}$ in the NS-PB and NS-KB settings. Using these choices of $\mathcal{A}_{\mathrm{e}}$, one can guarantee order-optimal regret in certain cases, as shown in the next section. The proofs of the following statements are given in the Appendix.
@@
Consider the NS-PB setting. Then, $\mathcal{A}_{\mathrm{e}}$ can be chosen as any arbitrary maximal linearly independent subset of $\mathcal{A}$
@@
Consider the NS-KB setting and assume w.l.o.g. that $\mathcal{A}\subseteq[0,R]^{d}$. Suppose that there exists an $\tilde{a}\in\mathcal{A}$ such that
@@
where $BL_{u}$ is the Lipschitz constant of all $f\in H_{k}(\mathcal{A})$. Let $\mathcal{P}_{T}\subset\mathcal{A}$ correspond to the set of the centers of the balls of an arbitrary $\delta_{T}$-cover. Then, the covering set $\mathcal{A}_{\mathrm{e}}$ can be taken as $\mathcal{P}_{T}$, with $\lvert\mathcal{P}_{T}\rvert\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
@@ 4.2 On Order-Optimality in Piecewise Stationary Environments
In the PS setting, the minimax regret lower bound under bandit feedback is $\tilde{\Omega}(\sqrt{N_{T}T})$ [24], [^3] We use the $\sim$ in $\tilde{\Omega}(\cdot)$ to hide polylogarithmic factors. which applies across all settings considered in this work, differing only in problem-dependent constants. Under certain conditions on the minimum spacing between change-points stated in [27], our algorithm matches this bound with state-of-the-art dependence on these constants. Specifically, the assumption states that $\nu_{k}-\nu_{k-1}$ should be large enough to trigger the change detector and timely restart.
Let $\ell_{\mathcal{D}}$ and $m_{\mathcal{D}}$ be the latency and the pre-change window length of the detector $\mathcal{D}$ [27], respectively. To characterize DAL’s performance under piecewise stationarity, we employ the methodology of [27], incorporating the regret analysis of the stationary bandit algorithm and that of the change detector. Theorem 4.3 easily follows from Theorem 1 from [27] and a proof sketch is provided in the Appendix.
@@
For the PS setting of Section 1, consider Algorithm 1 using the GLR test with parameters $\delta_{\mathrm{F}}$ and $\delta_{\mathrm{D}}$, a stationary bandit algorithm $\mathcal{B}$ with regret upper bound $R_{\mathcal{B}}$, a covering set $\mathcal{A}_{\mathrm{e}}$ and forced exploration frequencies $(\alpha_{k})_{k=1}^{\infty}$. Assume $\nu_{1}\geq\lceil N_{\mathrm{e}}/\alpha_{k}\rceil m_{\mathcal{D}}$ and $\nu_{k}-\nu_{k-1}\geq\lceil N_{\mathrm{e}}/\alpha_{k-1}\rceil\ell_{\mathcal{D}}+\lceil N_{\mathrm{e}}/\alpha_{k}\rceil m_{\mathcal{D}}$ for $k\in[2,N_{T}]$. If $R_{\mathcal{B}}(T)=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{T})$ with $p,q\geq 0$, $\delta_{\mathrm{F}}=\delta_{\mathrm{D}}=T^{-\gamma}$ for $\gamma>1$, and $\alpha_{k}=\alpha_{0}\sqrt{kN_{\mathrm{e}}\log\left(T\right)/T}$ with $a_{0}>0$, then DAL’s regret satisfies, $R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T}+\sqrt{N_{\mathrm{e}}N_{T}T\log T})$.
Using Theorem 4.3 and Propositions 4.1 and 4.2, we now provide the optimal regret DAL attains.
@@
Assume that the conditions of Theorem 4.3 hold. In the PS-PB setting, selecting $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.1, and in the PS-KB setting, selecting $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.2 with $\delta_{T}:=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ for some $C>0$, guarantees that DAL attains a regret
$R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}\sqrt{N_{T}T}).$
If the base stationary algorithm has order-optimal regret, DAL retains optimality in the PS setting. This also holds when $N_{\mathrm{e}}<d$ or $|\mathcal{A}|<d$ in PS-PBs, and when $N_{\mathrm{e}}<\gamma_{T}$ or $|\mathcal{A}|<\gamma_{T}$ in PS-KBs.

### v4

@@ 3.1 Synthetic Experiments
Figure 1: Dynamic regret vs. environment steps for synthetic experiments (lower==better). First three rows correspond to the geometric change-points and the final one to the drifting case.
@@ 3.1.1 Experimental Parameters
In all synthetic experiments, the action space comprises $100$ unique actions with dimension $d=10$. These actions are sampled independently from $\mathcal{N}(0,I)$. The horizon is fixed to $T=50000$ and we average the results over $15$ independent trials.
@@
In Alg. 1, when $|\mathcal{A}|$ is finite, change-detection can be performed on the actions selected by $\mathcal{B}$ that are not in $\mathcal{A}_{\mathrm{e}}$, which improves performance. This does not affect the theoretical properties of the algorithm, and we employ this variation for our experiments.
@@
The actions are scaled to lie within an $L$-ball and the underlying parameters $\theta_{t}$ belong to an $S$-ball. Specifically, for NS-LBs and NS-GLBs we have that $S=L=1$, while for NS-SCBs, we have that $L=1$ but $S=3$. Every time a $\theta_{t}$ is initialized or changed, its elements are chosen independently and uniformly in [1,1][-1,1], and then are scaled to the $S$-ball. For both NS-GLBs and NS-SCBs, we select $\mu(x):=\sigma(x)=(1+e^{-x})^{-1}$ (sigmoid). The additive noise $\varepsilon_{t}$ is sampled according to $\mathcal{N}(0,0.01)$ at each time-step, while for NS-SCBs, we sample the random reward according to $\mathrm{Bernoulli}(\mu(\langle\theta_{t},A_{t}\rangle))$ at time $t$. To set $\mathcal{A}_{\mathrm{e}}$ in NS-PBs, we use Corollary 4.5.
@@
Actions are scaled in the $\sqrt{d}$-ball and $\varepsilon_{t}\sim\mathcal{N}(0,0.01)$. Regarding the choice of kernel for NS-KBs, we employ the SE kernel with $\ell=0.2$. We follow a procedure similar to [17, 19]. Specifically, every time we initialize or change the reward function, $f_{t}$ is generated from the RKHS obtained by a discretization of [1,1][-1,1] into 200 evenly spaced points. The reward function is generated as $f(\cdot)=\sum_{i=1}^{M}\alpha_{i}k(\cdot,x_{i})$ with $\alpha_{i}$ uniformly distributed in [1,1][-1,1] and $M=200$. To identify $\mathcal{A}_{\mathrm{e}}$, we use Corollary 4.5.
NS-CBs
The context $C_{t}\in\mathbb{R}^{d_{c}}$ is drawn at each round from a fixed pool of 1000 normalized vectors with $d_{c}=10$, according to a categorical distribution. At every initialization or change, at least one of the $f_{t}$ or $\mathcal{P}_{t}$ changes. For $a\in\mathcal{A}$ and context $C_{t}$, $f_{t}$ is clipped in $[0,1]$, and is given by
$f_{t}(C_{t},a)=\Bigl[b_{a}+z^{(\mathrm{sig})}\,\sigma(u_{a}^{\top}C_{t})+z^{(\mathrm{sin})}\,\sin(v_{a}^{\top}C_{t})+z^{(\mathrm{xpr})}\,C_{t,2}C_{t,3}\Bigr]_{[0,1]},$
where $u_{a},v_{a}\sim\mathcal{N}(0,I)$, $b_{a}\sim\mathrm{Unif}[0.3,0.7]$, and $z^{\mathrm{(sig)}},z^{\mathrm{(sin)}},z^{\mathrm{(xpr)}}$ are drawn uniformly from $[0.25,0.45]$, $[0.15,0.35]$, $[0.10,0.25]$, respectively. Rewards are sampled as $\mathrm{Bernoulli}(f_{t}(C_{t},A_{t}))$. Since the reward function lacks any arm-related structure, here we set $\mathcal{A}=\mathcal{A}_{\mathrm{e}}$ (see Remark 4.3).
@@ 3.1.2 Experimental Benchmarks
In the PS setting, we adopt the geometric change-point model proposed in [28], and independently sample the intervals between the change-points according to a geometric distribution with parameter $\rho=T^{-\xi}$, for $\xi\in\{0.4,0.6,0.8\}$. We do not impose any restriction on the lengths of the intervals between change-points in our experiments.
@@
Regarding comparisons in drifting non-stationarity, we adopt the following drift model: in each run, the reward structure changes linearly over $T$ rounds from an initial value to a final value, where the end-points are chosen as in the beginning of the section. Specifically,
PBs: $\theta_{t}=(1-t/T)\,\theta_{\mathrm{init}}+(t/T)\,\theta_{\mathrm{final}},\quad\text{KBs:}\>\>f_{t}=(1-t/T)\,f_{\mathrm{init}}+(t/T)\,f_{\mathrm{final}},$
CBs: $\phi_{t}=(1-t/T)\,\phi_{\mathrm{init}}+(t/T)\,\phi_{\mathrm{final}},\>\phi_{t}:=(u_{a,t},v_{a,t},b_{a,t},\mathbf{z}_{t}),\ \mathbf{z}_{t}:=(z^{(\mathrm{sig})}_{t},z^{(\mathrm{sin})}_{t},z^{(\mathrm{xpr})}_{t})$.
@@
Per the results in Figure 1, DAL outperforms the current state-of-art methods in every synthetic experiment. DAL only abandons the actions chosen by the stationary bandit algorithm and restarts learning when an efficient change detector flags a mean-shift in rewards; hence, it avoids unnecessary restarts, especially when the intervals between the change-points are long enough for such detectors to correctly flag said changes without false alarms. Regarding drifting non-stationarity, DAL significantly outperforms all other methods. In fact, it fares better than both WeightUCB and ADA-OPKB, which not only are known to attain the optimal regret bound in the drift setup, but have also been shown to perform well in practice.
@@ 3.2 Real-World Experiments
Figure 2: Results for real-world experiments of Section 3.2, averaged over 15 independent runs. Top: dynamic regret (lower==better); Bottom: cumulative reward (higher==better).
Microarchitecture Prefetcher Selection Benchmark.
We introduce a novel dataset for NS bandit evaluation using the data of [29]. The dataset is derived from the SPEC06/17 benchmark suites, the gold standard in evaluating computer microarchitectural models. The dataset includes 11 prefetcher configurations (actions) that trade aggressiveness against efficiency. At each time-step, the reward is the normalized instructions per cycle in $[0,1]$, and the horizon is $T=26224$. We obtained the data directly from the authors. [^1] We aim to release the dataset to facilitate real-world experimentation by the bandit research community. Following [29], we also evaluate D-UCB [33] in its native form; while for our baselines we model the task as an NS-SCB. For reproducibility, D-UCB hyperparameters follow its original paper and [29]. Evaluation is by cumulative reward.
@@
NS-KBs have been applied to stock market prediction, and we follow the procedure of [19] to simulate two environments: one using their original data (Quandl stocks) and one constructed from NASDAQ-100 stocks retrieved via the yfinance Python package. [^2] Data retrieved from Yahoo Finance using the publicly available yfinance package. Used solely for non-commercial, academic research purposes. In the Yahoo-based dataset, we retain stocks with sufficient history (T=2000T{=}2000, approx. 5.5 years) and select the $50$ most volatile as actions. Daily closing prices define the reward function, and the empirical price covariance matrix is used as the kernel. To increase difficulty, we add Gaussian noise $\mathcal{N}(0,0.01)$ to the reward at each time-step. Evaluation is by dynamic regret.
COVID-NMA Clinical Benchmark.
We construct an NS-SCB benchmark from the open COVID-NMA database [10]. To maximize coverage while retaining clinical meaning, we form a UNION endpoint: for each bucketed-treatment arm and month, we include both Clinical Improvement at Day 28 (when reported) and Survival at Day 28 (1–mortality) as separate contributions, leading to binary rewards (1=success). Treatments (actions) are mapped into $13$ classes and month counts are expanded exactly ($s$ successes and nsn{-}s failures per bin) and concatenated in a fixed chronological order (month, clinD28 then mortD28, then bucket) to yield a long non-stationary sequence with $T\approx 7.4{\times}10^{4}$. Evaluation is based on cumulative reward.
Click Log Benchmarks.
We use the Yahoo! R6A click log dataset. [^3] Yahoo! Front Page Today Module User Click Log Datasets: https://webscope.sandbox.yahoo.com. Following prior works [12, 55], we compute smoothed click-through rates (CTRs) via rolling averages over 2000 rounds, average CTRs within each subperiod, and suppress fluctuations below $0.005$. To increase difficulty, we combine actions across 5 days, leading to 64 actions, compress the horizon to $50000$, and multiply final CTRs by 10. We model the resulting environment as an NS-SCB problem, reflecting the logistic reward structure typical in such settings [49]. Evaluation is by dynamic regret.
Alongside the first benchmark, we build a fixed-arm replay benchmark from the additional Yahoo! R6B click logs and cast it as an NS-CB problem.3 We select the highest-CTR articles to form an action set of $51$ actions. To improve coverage at a fixed horizon, we round-robin interleave days and then select $T=50000$. For each visit, we intersect the candidate set with this vocabulary, keep rounds where the displayed item remains, and record the binary click as the raw reward $(X_{t}\in\{0,1\})$. We rely on R6B’s uniform-random logging for unbiased replay/IPS evaluation [41]. Our metric is (replay) cumulative reward.
@@
We construct an NS bandit environment based on the Criteo live traffic dataset [20], following the preprocessing approach of [50] but modeling the problem as an NS-GLB rather than an NS-LB. We estimate the underlying parameter $\theta^{*}$ using logistic regression. Unlike [50], in which the authors employ a single change, we introduce shifts in $\theta^{*}$ via a geometric change-point model with parameter $\xi=0.8$ and extend the horizon to $T=50000$. The metric here is the dynamic regret.
@@
We use the Bioliq dataset provided by [34], which contains a week of measurements from 20 sensors in a power plant. We process the reward as [34] and construct an NS-SCB environment with $190$ actions. At each time-step, the reward is $1$ if the last 1000 measurements exceed a threshold of $2.04$, and 0 otherwise. Evaluation is based on cumulative reward.
@@ 4.1 On Effective Detection
When selecting a non-stationarity detector, accuracy and efficiency are essential for ensuring optimal regret growth. Any detector aiming to identify distribution shifts inherently requires a certain number of samples, both before and after the change. Ideally, this sample complexity should scale appropriately to avoid negatively impacting the total regret. To this end, GLR tests have been shown to achieve a pre- and post-change sample complexity of the order $\log T$ [32]. Since logarithmic terms are disregarded in dynamic regret analyses, it suggests that integrating this detection mechanism can achieve optimal regret growth.
To select which samples should be fed into the detector, one needs to properly select the covering set $\mathcal{A}_{\mathrm{e}}$, so that it contains actions that can capture changes in the reward function for any context. However, changes cannot be arbitrarily small, as no change detector may be able to identify them. Hence, $\mathcal{A}_{\mathrm{e}}$ should be designed such that whenever a change occurs, reward sequences associated with at least one context-action pair in $\mathcal{C}\times\mathcal{A}_{\mathrm{e}}$ exhibit an appreciable mean-shift. Define
$\Delta_{\mathrm{c}}\coloneqq\inf_{f\neq f^{\prime}}\max_{(c,a)\in\mathcal{C}\times\mathcal{A}_{\mathrm{e}}}\ \lvert f(c,a)-f^{\prime}(c,a)\rvert$.
Then, $\Delta_{\mathrm{c}}$ captures how well the context-action pairs in $\mathcal{C}\times\mathcal{A}_{\mathrm{e}}$ can discern between candidate reward functions. According to [31], $\Delta_{\mathrm{c}}$ crucially affects the performance of the GLR test, as its pre- and post- change sample complexity grows with $1/\Delta_{\mathrm{c}}^{2}$. The more discernible the changes are, the easier the detection becomes. Since forced exploration incurs regret, $\mathcal{A}_{\mathrm{e}}$ should ideally be chosen to minimize $N_{\mathrm{e}}$ while maximizing $\Delta_{\mathrm{c}}$. However, this cannot be done since the function $f_{t}$ is unknown. Hence, we provide the ways with which one can ensure appreciable mean-shift (i.e., $\Delta_{\mathrm{c}}>0$) in settings where the reward function has a certain structure (e.g., linear dependence on the arms or prescribed smoothness). Specifically, the NS-PB and NS-KB settings satisfy such conditions. Using these choices of $\mathcal{A}_{\mathrm{e}}$, one can guarantee order-optimal regret in certain cases, as shown in the next section. The proofs of the following propositions are given in the Appendix.
@@
In NS-PBs, $\mathcal{A}_{\mathrm{e}}$ can be any arbitrary maximal linearly independent subset of $\mathcal{A}$.
@@
In NS-KBs, assume that $\mathcal{A}\subseteq[0,R]^{d}$ w.l.o.g., and that there exists an $\tilde{a}\in\mathcal{A}$ s.t.
@@
where $BL_{u}$ is the Lipschitz constant of all $f\in H_{k}(\mathcal{A})$. Let $\mathcal{V}_{T}\subset\mathcal{A}$ be the set of the centers of the balls of an arbitrary $\delta_{T}$-cover. Then, $\mathcal{A}_{\mathrm{e}}$ can be taken as $\mathcal{V}_{T}$, with $\lvert\mathcal{V}_{T}\rvert\leq\lceil\sqrt{d}R/2\delta_{T}\rceil^{d}$.
Remark 4.3.
In NS-CBs, if $f_{t}$ and $\mathcal{A}$ satisfy the structural assumptions of the preceding propositions for any fixed context, we can set $\mathcal{A}_{\mathrm{e}}$ similarly. Without such structure, we set $\mathcal{A}_{\mathrm{e}}=\mathcal{A}$, as $\mathcal{A}$ is finite.
@@ 4.2 On Order-Optimality in Piecewise Stationary Environments
[…]
Let $\ell_{\mathcal{D}}$ and $m_{\mathcal{D}}$ be the latency and the pre-change window length of the detector $\mathcal{D}$ [31], respectively. To characterize DAL’s performance under piecewise stationarity, we employ the methodology of [31], incorporating the regret analysis of the stationary bandit algorithm and that of the change detector. Theorem 4.4 is based on Theorem 1 from [31], however, since we are studying general bandits, additional novel analysis is required. Thus, a proof sketch and the new analysis are provided in the Appendix.
@@
For the PS setting in Section 1, consider Alg. 1 using the GLR test with parameters $\delta_{\mathrm{F}}$ and $\delta_{\mathrm{D}}$, a stationary bandit algorithm $\mathcal{B}$ with regret upper bound $R_{\mathcal{B}}$, a covering set $\mathcal{A}_{\mathrm{e}}$ and forced exploration frequencies $(\alpha_{k})_{k=1}^{T}$. Assume $\nu_{1}$ $\geq$ $m_{1}$ and $\nu_{k}$ - $\nu_{k-1}$ $\geq$ $\ell_{k-1}$ ++ $m_{k}$ for $k$ $\in$ $\{2,\dots,N_{T}\}$, where $m_{k}=\lceil N_{\mathrm{e}}/\alpha_{k}\rceil m_{\mathcal{D}}$ and $\ell_{k}=\lceil N_{\mathrm{e}}/\alpha_{k}\rceil\ell_{\mathcal{D}}$ for $k\in[N_{T}]$ in PS-PBs and PS-KBs, and $m_{k}=\lceil N_{\mathrm{e}}/\alpha_{k}\rceil[\log(T)+m_{\mathcal{D}}/s]$ and $\ell_{k}=\lceil N_{\mathrm{e}}/\alpha_{k}\rceil[\log(T)+\ell_{\mathcal{D}}/s]$ for $k\in[N_{T}]$ with $s=\min_{c\in\mathcal{C},t\in[T]:\mathcal{P}_{t}(c)>0}\mathcal{P}_{t}(c)$ in PS-CBs. If $R_{\mathcal{B}}(T)=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(|\mathcal{A}|\log|\Pi|)^{r}\sqrt{T})$ with $p,q,r\geq 0$, $\delta_{\mathrm{F}}=\delta_{\mathrm{D}}=T^{-\gamma}$ for $\gamma>1$, and $\alpha_{k}=\sqrt{k|\mathcal{C}|N_{\mathrm{e}}}/(2\sqrt{T}\log^{2}T)$, then DAL’s regret satisfies, $R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(|\mathcal{A}|\log|\Pi|)^{r}\sqrt{N_{T}T}+\sqrt{|\mathcal{C}|N_{\mathrm{e}}N_{T}T})$.
In PS-PBs and PS-KBs, since the context is fixed, the minimum change-point separation $\ell_{k-1}+m_{k}$ is smaller. However, in PS-CBs, a context may appear with a low probability, and thus, the minimum change-point separation requires an extra $\log T$ term and an extra $s$ factor. Using Theorem 4.4 and Propositions 4.1 and 4.2, we now present the optimal regret DAL attains.
@@
Assume that the conditions of Theorem 4.4 hold. In PS-PBs, select $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.1. In PS-KBs, select $\mathcal{A}_{\mathrm{e}}$ as in Proposition 4.2 with $\delta_{T}:=\frac{Rd^{1/2-2p/d}}{2(C\gamma_{T}^{2q})^{1/d}}$ for some $C>0$. In PS-CBs, set $\mathcal{A}_{\mathrm{e}}$ as in Remark 4.3. Then, DAL attains
$R_{T}=\tilde{\mathcal{O}}(d^{p}\gamma_{T}^{q}(|\mathcal{A}|\log|\Pi|)^{r}\sqrt{N_{T}T}).$
If the base stationary algorithm has order-optimal regret, DAL retains optimality in PS-PBs, PS-KBs and PS-CBs. This also holds when $N_{\mathrm{e}}<d$ or $|\mathcal{A}|<d$ in PS-PBs, when $N_{\mathrm{e}}<\gamma_{T}$ or $|\mathcal{A}|<\gamma_{T}$ in PS-KBs, and when $\Pi$ is the universal set of all mappings from $\mathcal{C}$ to $\mathcal{A}$.